### v2

# Many-Worlds Inverse Rendering

Abstract.
Discontinuous visibility changes remain a major bottleneck when optimizing surfaces within a physically-based inverse renderer. Many previous works have proposed sophisticated algorithms and data structures to sample visibility silhouettes more efficiently.
Our work presents another solution: instead of differentiating a tentative surface locally, we differentiate a volumetric perturbation of a surface. We refer to this as a “many-worlds“ representation because it models a non-interacting superposition of conflicting explanations (“worlds”) of the input dataset. Each world is optically isolated from others, leading to a new transport law that distinguishes our method from prior work based on exponential random media.
The resulting Monte Carlo algorithm is simpler and more efficient than prior methods. We demonstrate that our method promotes rapid convergence, both in terms of the total iteration count and the cost per iteration.
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXXconference: ; ; journal: TOG
1. Introduction
Dramatic progress in the area of inverse rendering has led to methods that can fully reverse the process of image formation to reconstruct 3D scenes from 2D images.
This is usually formulated as an optimization problem: given a loss function $\mathcal{L}$ and a rendering $\mathcal{R}(\pi)$ of tentative parameters $\pi$, we seek to minimize their composition
(1) $\pi^{*}=\operatorname*{argmin}_{\pi\in\Pi}\mathcal{L}(\mathcal{R}(\pi)).$
The details greatly vary depending on the application, but usually $\mathcal{L}$ will quantify the difference between the rendering and one or more reference images.
Most recent work on this problem builds on the notion of emissive volumes [Mildenhall et al., 2021; Kerbl et al., 2023]. They are popular because the direct relationship between emission and observations leads to exceptionally well-behaved optimization problems.
Physically based methods instead seek a more complicated explanation: they simulate emission and scattering inside a general scene that can contain essentially anything: surfaces, fibers, volumes, physically-based BSDFs, etc. Interreflection adds dense nonlinear parameter dependencies that make this a significantly harder optimization problem. It goes without saying that these two approaches do not compete: when an emissive volume is an acceptable answer, it should always be preferred. This is because it encapsulates material, lighting, geometry and inter-reflection in a unified field representation that promotes speedy and stable convergence.
That said, many applications besides novel view synthesis rely on physically interpretable reconstructions. Unfortunately, algorithms designed to address this harder problem are often surprisingly brittle, which is especially apparent when these methods are tasked to reconstruct 3D shapes. Our work extends previous theory and opens up a new direction to confront this instability.
The default strategy to solve Equation 1 is gradient-based descent:
(2) $\vspace{-.5mm}\pi_{i+1}\leftarrow\pi_{i}-\eta\nabla_{\pi}\mathcal{L}(\mathcal{R}(\pi_{i})),\vspace{-.5mm}$
involving various refinements (moments, regularization, etc.) that stabilize and accelerate convergence.
In this process, we often already have a particular kind of solution in mind (e.g., a triangle mesh), which constrains the domain $\Pi$ containing $\pi^{*}$. It seems only natural that we would also try to reach this goal by evolving compatible models $\pi_{i}\in\Pi$ along the way, e.g., by deforming a tentative surface starting from an initial guess.
However, optimizing surfaces in this way is unreasonably hard. Computing $\nabla_{\pi}\mathcal{L}(\mathcal{R}(\pi_{i}))$ with automatic differentiation produces incorrect gradients due to parameter-dependent discontinuities that riddle the integration domain of $\mathcal{R}$. Incorporating specialized techniques to fix this problem simply leads to the next one: the algorithm now outputs sparse gradients on object silhouettes that cause convergence to bad local minima. Adding further precondition and regularization can help. However, even with all of these fixes in place, the optimization often still does not work all that well.
But does it need to be this way? Taking the Chinese proverb
There are many paths to the top of the mountain, but the view is always the same.
to heart, Equation 2 is merely a means to an end. To the extent that this simplifies the inversion, there is no fundamental reason why we cannot temporarily depart the set $\Pi$ to consider other alternative explanations, potentially even ones that are not readily explainable in terms of traditional physically based rendering.
We consider optimizations on an extended parameter space
[…]
where $\tilde{\mathcal{R}}$ is furthermore parameterized by $\tilde{\pi}\in\tilde{\Pi}$ representing features that we are unwilling to accept in the final solution. If the role of these “perpendicular dimensions” diminishes over time, i.e., if
[…]
then we can simply discard $\tilde{\pi}^{*}$ at the end and take $\pi^{*}$ to be the solution of the original optimization problem.
The parameter space extension serves two purposes: first, it turns a circuitous trajectory through a non-convex energy landscape into one that exploits the additional degrees of freedom to make more direct progress towards the solution:
Second, we will choose the extension so that visibility changes in the original problem cease to be discontinuous in the extended space, which greatly reduces the difficulty of gradient estimation.
Seen from a high level, this isn’t a new idea: for example, numerous prior works have observed how surface reconstruction problems become tractable once those surfaces have been “blurred” into a volume. A recent paper by Miller et al. [2024] interprets microflake volumes as the relaxation of a stochastic surface model and discusses a large body of related work that optimizes volumes to eventually obtain surfaces. While the extension from surfaces to exponential volumes does indeed address many problems, we argue that it is unnecessarily complex:
  1. (1)
    Speed. The presence of random volumes in a physically based renderer implies that it must solve the radiative transfer equation (RTE) requiring free-flight sampling and transmittance estimation along every ray segment. In heterogeneous volumes, both are costly iterative methods.
  2. (2)
    Shadowing. The RTE models the effect of randomly distributed particles within an infinitesimal cylindrical region surrounding a ray. Free-flight sampling and transmittance estimation have to be iterative because they must resolve how particles at one location along this ray could influence others by redirecting or absorbing a portion of the illumination.
    However, these particles aren’t real (or at least, not yet). They constitute possible explanations of surfaces that may or may not become part of the final reconstruction. It makes little sense for them to shadow each other.
  3. (3)
    Extraction. Turning fuzzy inter-reflecting volumes into optically equivalent flattened surfaces presents its own challenges. Prior work related the bulk properties of opaque microfacets and microflakes [Dupuy et al., 2016; Miller et al., 2024], but this insight does not generalize to the richness of reflectance models in use in today’s rendering systems. In general, a complete pipeline will require a post-process optimization step to find a surface that locally matches a previously reconstructed volume. This breaks the elegance of simply discarding extra dimensions once the optimization converges (Equation 4).
Instead of modeling a volume, in which the effects of particles manifest multiplicatively (leading to exponential volumes), we develop a model where a random distribution of surfaces manifests as an additive perturbation of the scene.
We refer to this as a “many-worlds” representation because it describes many non-interacting alternative explanations of the input data (“worlds”). Furthermore, instead of entirely replacing the surface with a volume during the optimization, our optimization keeps the surface and augments it with a spatial representation that models possible perturbations of that surface.
Differentiation of this model produces dense surface derivatives in an extended domain without explicit silhouette sampling, and the simpler non-interacting nature of the representation also reduces the cost of optimization steps compared to exponential volumes. Finally, mesh extraction is a simple projection that discards $\tilde{\pi}$, and all of this is easily incorporated into a physically based path tracer.
Our main contribution is a new solution to the classic discontinuity problem that is simpler and more efficient than prior methods. We present two derivations that alternatively formulate the central algorithm as either
  • a first-order volumetric perturbation of an evolving surface (Section 3), or
  • an extension of the surface derivative domain that removes the need for silhouette sampling (Section 4 and the Appendix).
Although this paper will occasionally use the word ”volume” to refer to the many-worlds perturbation of the scene, our method should not be understood as a volume rendering method in the sense of traditional media: it employs BRDFs rather than phase functions, lacks scattering coefficients, and has no concept of transmittance.
2. Related work and background
This section reviews related prior work on inverse rendering of solid geometry. The disentanglement of geometry, material and lighting is orthogonal to the topic of this work.
2.1. Physically-based inverse rendering
The physically-based approach tries to model all available information in input images by accounting for lighting, materials, and interreflection. This leads to several challenges.
Discontinuities
Analytic differentiation of the rendering equation produces a boundary derivative term that is not sampled by primal rendering strategies. Consequently, applying automatic differentiation to a rendering algorithm produces incorrect gradients.
Estimating the missing derivative requires sampling rays along shape boundaries (i.e., silhouettes). Prior work [Li et al., 2018; Zhang et al., 2020; Yan et al., 2022; Zhang et al., 2023a] has extensively studied the theory and practice of estimating this boundary term. Another approach employs reparameterizations to convert the boundary contribution to neighboring regions, enabling simpler area-based formulations. These methods [Loubet et al., 2019; Bangaru et al., 2020] often require tracing auxiliary rays to detect the presence of neighboring silhouettes. For geometry representations based on signed distance fields (SDFs), specialized algorithms can alleviate some of this cost [Vicini et al., 2022; Bangaru et al., 2022].
While the theory of the boundary term continues to evolve, practical efficiency and robustness remain challenging. Current methods still have important limitations, particularly concerning bias (e.g., specular surfaces) and variance (e.g., when multiple silhouettes are in close proximity). These methods are difficult to implement, and the cost of boundary handling is usually the main bottleneck of the entire optimization.
Sparse gradients
The derivative of the rendering process can be decomposed into two types: the continuous part and the discontinuous part. We refer readers to Figure 4 in [Zhang et al., 2023a] for an illustration. Since our paper focuses on geometry optimization, we refer to the continuous part as the shading derivative and the discontinuous part as the boundary derivative in the following.
The shading derivative is defined everywhere on the surface and primarily affects the surface through its normal. Because shading depends on lighting, the smoothness of the illumination environment can have a marked effect on the robustness of optimization. Adjoint rendering methods [Nimier-David et al., 2020; Vicini et al., 2021b] exploit symmetries to compute the shading derivative with linear time complexity. Zeltner et al. [2021] analyze different strategies for variance reduction in this process. The boundary derivative, on the other hand, is only defined on the visibility silhouette of the surface. These surface derivatives are sparse in two senses:
  1. (1)
    A single view only observes a limited subset of all possible silhouettes, and a derivative step is likely to introduce a kink at this subset. Nicolet et al. [2021] reparameterize meshes on a space that promotes smoothness to mitigate this issue.
  2. (2)
    The derivatives are only defined on the surface, not throughout the entire 3D space. Although light transport is simulated in this larger space, only the neighborhood of a surface can use the computed gradients. Regions far from the initial guess, must wait until the surface extends to them to receive gradients. Mehta et al. [2023] propose an elegant solution for vector graphics in 2D; however, they do not extend their approach to define surface derivatives in the entire 3D space.
The sparsity has two implications: (1) the optimization is slow because it needs many iterations to deform the surface to the target shape, and (2) a good initial guess is needed, especially when the target shape has a complicated topology or self-occlusions.
Approximations
There is a substantial body of literature on approximating the PBR process to make the optimization more tractable. These methods often bypass the discontinuity problem by assuming knowledge of a shape mask (implicitly assuming the shape is directly observable), or by using approximated boundary derivative estimation algorithms like Nvdiffrast [Laine et al., 2020]. Additionally, most methods approximate the rendering equation for efficiency [Zhang et al., 2021a; Jin et al., 2023; Zhang et al., 2021b; Boss et al., 2021]. These methods are physically-inspired and can handle complex scenes well, while our theory focuses on differentiating a fully physically-based rendering process.
Nimier-David et al. [2022] demonstrate that a physically-based non-emissive volume can be optimized to fit the input images. These methods are more physically accurate than the emissive volume methods but do not yield a surface representation in the end.
2.2. Emissive volume inverse rendering
A large body of work based on variants of the NeRF [Mildenhall et al., 2021] technique replaces the expensive rendering process with an emissive volume, which leads to convergent optimizations with excellent generalization to novel viewpoints. However, the final reconstruction is fog-like volume that does not accurately represent a solid surface.
Later work [Yariv et al., 2021; Wang et al., 2021; Li et al., 2023; Miller et al., 2024] added a surface prior to the volume model by introducing an SDF and deriving the volume density from it. Intuitively, volume particles should concentrate near the surface and decrease as one moves away from it. To model this effect, a symmetric bell-shaped distribution is added to the SDF value (centered at the zero-level set) to convert it into a volume density. The variance of the distribution is learnable during optimization: it begins with a high-variance distribution, mimicking the NeRF model, and gradually reduces to ensure that a surface can confidently be extracted in the end.
These SDF-based volume reconstructions can be seen as a blurry version of the surface-based methods without indirect effects. When the variance is zero, these methods simplify to surface-based methods [^1] This equivalence is theoretical. In practice, algorithms designed for volume optimization cannot handle such a Dirac-delta density distribution due to discontinuities.. With moderate variance, the effect of a solid surface, which we can interpret as “terminating a ray at a position with probability $1$”, is now diffused to the neighborhood, interpretable as “terminating a ray in this neighborhood according to a probability density”. The blur eliminates the need to compute boundary derivatives.
A general limitation of all methods in this category is that they fail to accurately account for global effects like shadowing and interreflection.
2.3. Surfaces versus volumes
Several related works studied the relationship of surface and volume rendering: Heitz et al. [2016] and Dupuy et al. [2016] describe microfacet surfaces within the framework of random volumes. Vicini et al. [2021a] incorporated non-exponential transmittance to account for correlations arising from opaque surfaces. Miller et al. [2024] formalize microflake volumes as the relaxation of a stochastic surface model and solve inverse problems using this representation. Seyb et al. [2024] recently incorporated a richer set of spatial correlations, producing a forward rendering framework that spans the full continuum ranging from pure volume to surface-like interactions.
3. Method
We first begin with an informal motivation of our method followed by more precise definitions. The appendix provides an alternative derivation of the central equation starting from prior work on physically-based differentiable rendering.
3.1. Motivation
Consider the notion of a stochastic surface $S$ defined by a process that randomly generates surfaces according to a probability distribution. We shall think of a single sample from this distribution as a complete 3D scene representing one of many possible “explanations” of the optimization’s objective and will refer to it as a world.
Miller et al. [2024] recently studied the relationship between random surface distributions and microflake volumes. They showed that under reasonable assumptions, the distribution of surfaces can be aggregated into a 3D occupancy
(5) $\alpha(\mathbf{x})\coloneqq\Pr\{\mathbf{x}\text{ is inside of }S\},$
and that this occupancy $\alpha$ can serve as the foundation of an exponential microflake volume $V_{S}$, where it leads to a sensible (energy-conserving, reciprocal) light transport model.
Such a volumetric rendering simulates a kind of average across different worlds, each weighted by its associated probability density. However, the details of this averaging process are difficult to reason about because averaging surfaces into a volume leads to interactions that bridge multiple worlds. In particular, the rendering of the averaged world is not the average of renderings of the individual worlds:
(6) $\mathcal{R}(V_{S})\neq E[\mathcal{R}(S)]$
where $\mathcal{R}$ is a generalized volume/surface renderer in this context.
To see why, consider a distribution $S$ describing a nearly empty scene containing a single infinite plane with an uncertain offset. Within any particular world, light traveling towards this plane will scatter exactly once. In contrast, the volume model $V_{S}$ blurs $S$ into a band containing a distribution of microflakes that cause arbitrarily long chains of scattering events. Not only is this expensive to simulate—it also discombobulates the optical interpretation of these interactions as the effect of a surface to be reconstructed.
We also optimize a distribution of worlds but treat them independently. We do this by changing parameters (e.g., material properties) within individual worlds, and by changing the relative probability with which these worlds occur. However, we never consider nonsensical inter-world scattering or shadowing.
Our method presumes the existence of a non-random mean surface $\bar{S}$, which is the expected value of the random surface $S$.
(7) $\bar{S}:=E[S]$
The original and extended parameter space of our optimization are then given by $\pi=\bar{S}$ and $\tilde{\pi}=S$. Here, it is important to already mention that we will never optimize the mean surface $\bar{S}$ directly. Yet, this surface still evolves during the optimization, which is a consequence of changes to the distribution $S$ that indirectly affect the mean. At the end of the optimization, we keep $\bar{S}$ and discard $S$.
A many-worlds rendering computes an additive volumetric perturbation of the surface rendering of $\bar{S}$:
(8) $\mathcal{R}(\pi,\tilde{\pi})=\mathcal{R}(\bar{S})+\lim_{\varepsilon\to 0}\frac{\mathcal{R}(\bar{S}+\varepsilon\cdot V_{S})-\mathcal{R}(\bar{S})}{\varepsilon},$
which has the following explanation:
  1. (1)
    We first render the surface $\bar{S}$, which produces a regular image. We then add a second component that represents other worlds.
  2. (2)
    The term $\mathcal{R}(\bar{S}+\varepsilon\cdot V_{S})$ in the numerator renders a scene containing both the mean surface and the volume $V_{S}$ whose density has, however, been diluted by a factor $\varepsilon$. The resulting image is thus nearly identical to $\mathcal{R}(\bar{S})$ since the medium is so thin.
  3. (3)
    We compute the difference between these two images, divide by $\varepsilon$, and let $\varepsilon\to 0$. In other words, we take a derivative.
This expression can be understood as a first-order Taylor expansion of a surface rendering evaluated using a volumetric perturbation. While this formulation is of little use for traditional rendering, we will demonstrate its effectiveness once differentiated. Said differently, many-worlds rendering is primarily designed as a framework for derivative propagation rather than radiance propagation.
Equation 8 explains what many-worlds rendering mathematically models, but the actual definition in terms of a limiting process of a rendering algorithm is not exactly straightforward. Our next goals are therefore to clarify the implications of this equation, and to convert it into a form that can be implemented in practice.
To structure this explanation in terms of well-established ideas, we go through an at times lengthy sequence of steps.
  • Section 3.2 reviews classic exponential random volumes.
  • Section 3.3 looks at how random volumes simplify when their density is very low.
  • Section 3.4 establishes the connection between surfaces and volumes by defining optical volume parameters representing surfaces. This builds on anisotropic radiative transport, which the section also reviews.
  • Section 3.5 switches back to a thick volume, while retaining the simplified properties of the low-density case.
  • Section 3.6 differentiates the final model to show how derivatives propagate in the many-worlds representation.
At the end, most of the intermediate complexity will simply evaporate, leaving behind a surprisingly simple model that can be implemented in a few lines of code.
Our method departs from prior work in the following notable way: whereas existing inverse rendering methods have largely differentiated known surface-based or volume-based rendering algorithms, our work proposes a dedicated derivative transport law that operates within the derivative propagation step of radiative backpropagation [Nimier-David et al., 2020]. This means that the resulting derivatives do not correspond to the differentiation of any physically-based method. Indeed, a direct attempt to retrofit them to primal rendering would produce nonsensical results that violate energy conversation. Section 3.7 finally revisits this point and shows how the many-worlds representation can be adapted to the setting of primal rendering.
3.2. Transport in an exponential random volume
We begin by reviewing light transport in an exponential random volume as described by the integral form of the radiative transfer equation (RTE). This equation gives the incident radiance $L_{\mathrm{i}}$ along a ray $(\mathbf{x},\bm{\omega})$, accounting for radiative gains (in-scattering, emission) and losses (out-scattering, absorption) along the segment reaching up to the nearest ray-surface intersection at $s=\inf\{s^{\prime}\mid\mathbf{x}+s^{\prime}\bm{\omega}\in\mathcal{M}\}$ or $s=\infty$ if none exists:
(9) $L_{\mathrm{i}}(0)=\int_{0}^{s}T(t)\left[\mu_{\mathrm{s}}(t)\,L_{\mathrm{s}}(t)+\mu_{\mathrm{a}}(t)\,L_{\mathrm{e}}(t)\right]\mathrm{d}t+T(s)\,L_{\mathrm{o}}(s),$
(where we have taken notational liberty by only parameterizing quantities based on the ray distance $t$ to keep things simple). The first term models contributions of the volume, while the second accounts for a potential surface at $t=s$, whose outgoing radiance $L_{\mathrm{o}}(s)$ is attenuated by the medium.
The functions $\mu_{\mathrm{a}}(t)$ and $\mu_{\mathrm{s}}(t)$ specify the volume’s absorption and scattering coefficient, whose sum gives the extinction $\mu_{\mathrm{t}}=\mu_{\mathrm{a}}+\mu_{\mathrm{s}}$. The in-scattered radiance
(10) $L_{\mathrm{s}}(t)=\int_{\mathcal{S}^{2}}L_{i}(t,\bm{\omega}^{\prime})\,f_{p}(t,\bm{\omega},\bm{\omega}^{\prime})\,\mathrm{d}\bm{\omega}^{\prime}$
is the product integral of incident radiance and the phase function $f_{p}$. Finally, the transmittance
(11) $T(t)=\exp\left(-\int_{0}^{t}\mu_{t}(t^{\prime})~{}\mathrm{d}t^{\prime}\right)$
ties everything together: it establishes the connection to particle distributions and ensures energy conservation by accounting for self-shadowing (or more generally, self-extinction, see Figure 2a).
3.3. Transport in a tenuous volume
A many-worlds rendering describes a perturbation by a tenuous volume: its density in Equation 8 is diluted by a factor of $\varepsilon$, which reduces the scattering and extinction coefficients $\mu_{\mathrm{s}}$ and $\mu_{\mathrm{t}}$. This low amount of extinction enables an approximation of Equation 11:
(12) $T(t)\approx 1-\int_{0}^{t}\mu_{t}(t^{\prime})\,\mathrm{d}t^{\prime},$
which follows from $e^{-x}\approx 1-x+O(x^{2})$. Plugging this into the RTE (9) and discarding a second-order term ($\mu_{\mathrm{s}}\mu_{\mathrm{t}}\sim\varepsilon^{2}$) yields the approximation $L_{\mathrm{i}}(0)\approx L_{\mathrm{i}}^{\varepsilon}(0)$, where the latter is defined as
(13) $L_{\mathrm{i}}^{\varepsilon}(0)\!\coloneqq\!\!\!\int_{0}^{s}\!\!\!\left[\mu_{\mathrm{s}}(t)\,L_{\mathrm{s}}(t)+\mu_{\mathrm{a}}(t)\,L_{\mathrm{e}}(t)-\mu_{\mathrm{t}}(t)\,L_{\mathrm{o}}(s)\right]\mathrm{d}t+L_{o}(s),\!$
In other words, a tenuous volume is governed by a linearized RTE, where higher-order terms vanish. In the following, variables with a superscript $\varepsilon$ refer to quantities in such a linearized medium with dilution factor $\varepsilon$.
Figure 2. (a) Radiance $L_{o}$ passing through an exponential random medium undergoes extinction due to a distribution of particles. Their effect compounds, producing a nonlinear shadowing effect. (b) The many-worlds approach models an additive superposition of different surfaces. Light can never interact with multiple worlds—consequently, they cannot shadow each other.
3.4. Completing the volume model
To complete this model, we must instill meaning into the terms $\mu_{\mathrm{t}}$, $\mu_{\mathrm{s}}$, and $f_{p}$. Prior work often did so using volumetric analogs of microfacet surface models known as microflakes [Jakob et al., 2010]. Applications include optimization, volumetric level of detail, energy-conserving random walks, and neural fields [Heitz et al., 2015, 2016; Vicini et al., 2021a; Loubet and Neyret, 2018; Zhang et al., 2023b].
It is worth noting that the theory of microflakes actually builds on a more general notion of anisotropic radiative transport [Shultis and Myneni, 1988; Jakob et al., 2010], which adopts directionally varying extinction and scattering coefficients arising from a random distribution of oriented particles:
(14) $\mu_{\mathrm{t}}(\mathbf{x},\bm{\omega})$ $=\rho(\mathbf{x})\int_{S^{2}}\sigma(\bm{\omega},\bm{\omega}^{\prime})\,D(\mathbf{x},\bm{\omega}^{\prime})\,\mathrm{d}\bm{\omega}^{\prime}$
(15) $\mu_{\mathrm{s}}(\mathbf{x},\bm{\omega})$ $=\rho(\mathbf{x})\int_{S^{2}}a(\mathbf{x},\bm{\omega},\bm{\omega}^{\prime})\,\sigma(\bm{\omega},\bm{\omega}^{\prime})\,D(\mathbf{x},\bm{\omega}^{\prime})\,\mathrm{d}\bm{\omega}^{\prime}$
where $\rho(\mathbf{x})$ is the particles’ number density at $\mathbf{x}$, $D(\bm{\omega})$ models the density of their directional orientations, $\sigma(\bm{\omega},\bm{\omega}^{\prime})$ gives the cross-sectional area of a single particle with orientation $\bm{\omega}^{\prime}$ observed from direction $\bm{\omega}$, and $a(\mathbf{x},\bm{\omega},\bm{\omega}^{\prime})$ models the particles’ scattering albedo ($\in[0,1]$). Microflake theory can then be derived from these expressions by setting $\sigma(\bm{\omega},\bm{\omega}^{\prime})=\sigma\,|\bm{\omega}\cdot\bm{\omega}^{\prime}|$ to model the projected area of a facet with surface area $\sigma$, and by constructing a phase function $f_{p}$ based on the principle of specular reflection from such a flake.
However, we do not use microflakes in this work. Instead, we adopt a far simpler particle model that associates a single particle orientation $\bm{\beta}(\mathbf{x})$ with every point $\mathbf{x}$ (Figure 3).
Figure 3. Occupancy and orientation fields. The above images visualize the contents of an occupancy ($\alpha$) and orientation ($\bm{\beta}$) field following an optimization. The former models the probability of a surface existing at a position, while the latter assigns normal directions.
Since there is only one orientation at $\mathbf{x}$, the distribution $D$ collapses to a Dirac delta function: $D(\mathbf{x},\bm{\omega})=\delta(\bm{\omega}-\bm{\beta}(\mathbf{x}))$, which in turn reduces the extinction to a simple product of number density and cross-sectional area.
(16) $\mu_{\mathrm{t}}(\mathbf{x},\bm{\omega})=\rho(\mathbf{x})\,\sigma(\bm{\omega},\bm{\beta}(\mathbf{x}))$
Here, we do not specifically model the split into a separate number density and cross-section. Instead, we define an extinction function that directly evaluates their product:
(17) $\mu_{\mathrm{t}}(\mathbf{x},\bm{\omega})$ =$=\begin{cases}\varepsilon\,\alpha(\mathbf{x}),&\mathrm{if}\ \bm{\omega}\cdot\bm{\beta}(x)>0,\\ 0,&\mathrm{otherwise},\end{cases}$
where $\varepsilon$ is the dilution factor and $\alpha(\mathbf{x})$ is the fractional occupancy, i.e., the point-wise discrete probability that $\mathbf{x}$ is inside the surface $S$ (Eq. 5). The branch condition encodes an important non-reciprocal [^2] This behavior is non-reciprocal because when the medium interacts along $\bm{\omega}$ (i.e., if $\mu_{t}(\mathbf{x},\bm{\omega})>0$), then the opposite direction is non-interacting (i.e., $\mu_{t}(\mathbf{x},-\bm{\omega})=0$). behavior: light interacting with a back-facing surface presents a nonsensical case, and this term masks those parts of the volume.
The extinction $\mu_{\mathrm{t}}$ models the stopping power of the volume. For example, the flakes of a microflake volume obstruct light to a lesser degree when it propagates nearly parallel to them, and this manifests via a foreshortening term $|\bm{\omega}\cdot\bm{\omega}^{\prime}|$ in the definition of $\mu_{\mathrm{t}}$. This is important to obtain a sensible (energy-conserving, reciprocal) model because these particles mutually interact. On the other hand, when focusing on a single world in a many-worlds representation, light stops with probability 1 when it encounters a surface regardless of how it is oriented (except for mentioned back-facing case). This is why our model does not include a similar foreshortening term. The term $\alpha(t)$ models the discrete probability of the world within the ensemble. While reciprocity is often an important concern in simulations, it plays a lesser role here since the volume merely perturbs an existing simulation that does satisfy reciprocity.
Figure 4. Our method differentiates a volumetric perturbation of a surface rendering to derive a novel derivative transport law. The scene contains a distribution of possible surfaces and its mean surface $\bar{S}$ (which may consist of many disconnected components). A light path can interact at most once with the stochastic many-worlds representation (shaded region). The quantities $\alpha(\mathbf{x})$ and $\bm{\beta}(\mathbf{x})$ define the behavior of possible surfaces in space. The occupancy $\alpha(\mathbf{x})$ gives the discrete probability that $\mathbf{x}$ is inside the surface, and $\bm{\beta}(\mathbf{x})$ provides a local surface orientation.
We define the phase function $f_{p}$ of the volume as
$f_{p}(\mathbf{x},\bm{\omega},\bm{\omega}^{\prime})=\frac{f_{s}(\mathbf{x},\bm{\omega},\bm{\omega}^{\prime})\,|\bm{\beta}(\mathbf{x})\cdot\bm{\omega}^{\prime}|}{a(\mathbf{x},\bm{\omega},\bm{\beta}(\mathbf{x}))},$
which wraps $f_{s}$, the bidirectional scattering distribution function (BSDF) of the many-worlds surface at the point $\mathbf{x}$. This definition follows the standard convention that the phase function integrates to $1$, with absorption handled by other terms. The albedo $a$ provides the necessary normalization constant:
(18) $a(\mathbf{x},\bm{\omega},\bm{\beta}(\mathbf{x}))=\int_{S^{2}}f_{s}(\mathbf{x},\bm{\omega},\bm{\omega}^{\prime})\,|\bm{\beta}(\mathbf{x})\cdot\bm{\omega}^{\prime}|\,\mathrm{d}\bm{\omega}^{\prime}.$
With these definitions, $\mu_{\mathrm{s}}$ reduces to the product of extinction and $a$:
(19) $\mu_{\mathrm{s}}(\mathbf{x},\bm{\omega})$ $=a(\mathbf{x},\bm{\omega},\bm{\beta}(\mathbf{x}))\,\mu_{\mathrm{t}}(\mathbf{x},\bm{\omega}).$
The product of this scattering coefficient and in-scattered radiance $L_{\mathrm{s}}$ (Equation 10) can be seen to compute an extinction-weighted integral of the many-worlds BSDF over projected solid angles.
$\mu_{\mathrm{s}}(\mathbf{x},\bm{\omega})L_{\mathrm{s}}(\mathbf{x},\bm{\omega})=\mu_{\mathrm{t}}(\mathbf{x},\bm{\omega})\!\!\int_{S^{2}}\!\!L_{\mathrm{i}}(\mathbf{x},\bm{\omega}^{\prime})\,f_{s}(\mathbf{x},\bm{\omega},\bm{\omega}^{\prime})\,|\beta(\mathbf{x})\cdot\bm{\omega}^{\prime}|\,\mathrm{d}\bm{\omega}^{\prime}$
We further use the sum of the above expression with the absorption-weighted emission to define an extended outgoing radiance function $L_{\mathrm{o}}$ for points $\mathbf{x}$ that lie within the many-worlds representation:
(20) $\mu_{\mathrm{s}}(\mathbf{x},\bm{\omega})L_{\mathrm{s}}(\mathbf{x},\bm{\omega})+\mu_{\mathrm{a}}(\mathbf{x},\bm{\omega})L_{\mathrm{o}}(\mathbf{x},\bm{\omega})$ $\eqqcolon\mu_{\mathrm{t}}(\mathbf{x},\bm{\omega})L_{\mathrm{o}}(\mathbf{x},\bm{\omega}).$
If we substitute all of these expressions into Equation 13, we obtain
(21) $L_{\mathrm{i}}^{\varepsilon}(0)=\varepsilon\int_{0}^{s}\alpha^{+}(t)\left[L_{\mathrm{o}}(t)-L_{\mathrm{o}}(s)\right]\mathrm{d}t+L_{o}(s),$
where $\alpha^{+}$ gives the occupancy at $t$ and is zero in the back-facing ($\bm{\beta}(t)\cdot\bm{\omega}<0$) case.
This equation clarifies how the many-worlds representation affects transport along one segment of a light path. However, we usually render longer paths with many segments. In this case, only one segment will use Equation 21. We do not permit multiple interactions with the representation, because this, again, would be a second-order effect proportional to $\varepsilon^{2}$ that vanishes for small $\varepsilon$ (Figure 4).
3.5. Many-worlds transport
Recall Equation 8, which introduced the idea of a volumetric perturbation of a surface rendering:
$\mathcal{R}(\pi,\tilde{\pi})=\mathcal{R}(\bar{S})+\lim_{\varepsilon\to 0}\frac{\mathcal{R}(\bar{S}+\varepsilon\cdot V_{S})-\mathcal{R}(\bar{S})}{\varepsilon}.$
The term $\mathcal{R}(\bar{S}+\varepsilon\cdot V_{S})$ in the numerator led to the derivation of the linearized transport law (21), which produces this perturbed image. The effect of subsequently subtracting $\mathcal{R}(\bar{S})$, dividing by $\varepsilon$, and then adding back $\mathcal{R}(\bar{S})$ is to cancel out the $\varepsilon$ scale factor so that the final transport law has a unit amount (thus $L_{\mathrm{i}}^{1}$) of the many-worlds perturbation:
(22) $L_{\mathrm{i}}^{1}(0)=\int_{0}^{s}\alpha^{+}(t)\left[L_{\mathrm{o}}(t)-L_{\mathrm{o}}(s)\right]\,\mathrm{d}t+L_{o}(s).$
This equation captures how occupancy $\alpha^{+}(t)$ at position $t$ adds a contribution from the many-worlds perturbation, while at the same time occluding a background surface at $s$, if present.
3.6. Many-worlds derivative transport
The many-worlds formulation exists to be differentiated. As mentioned before, the mean surface does not participate in this process, which means that some of the terms of Equation 22 disappear under differentiation with respect to a parameter $\pi$:
[…]
(23) $=\int_{0}^{s}\big{(}\underbracket{\partial_{\pi}\alpha^{+}(t)\left[L_{\mathrm{o}}(t)-L_{\mathrm{o}}(s)\right]}_{\text{(i)}}+\underbracket{\alpha^{+}(t)\partial_{\pi}L_{\mathrm{o}}(t)}_{\text{(ii)}}\big{)}\,\mathrm{d}t$
The equation shows that when the optimization tries to increase the incident radiance at $t=0$, it can do so in two ways:
  1. (i)
    By increasing the occupancy if $L_{\mathrm{o}}(t)>L_{\mathrm{o}}(s)$ or decreasing it otherwise.
  2. (ii)
    By increasing the outgoing radiance $L_{\mathrm{o}}(s)$, e.g., using a change to the $\bm{\beta}$, or other potential material parameters referenced by the BSDF (e.g., an albedo).
The only way in which the mean surface manifests at this step is to provide a baseline radiance value that enables computing the color difference in term (i).
Note that the derivative propagation described by Equation 23 lacks a transmittance term that would ordinarily model attenuation by the occupancy along the ray (Figure 5). The property stems from the principle that different worlds should not shadow each other.
This concludes our derivation of many-worlds derivative transport using the concept of first-order volumetric perturbation of an evolving surface. In Section 4 and the Appendix, we reveal the essence of this novel transport using the surface derivative theory developed by prior works.
Figure 5. Many-worlds derivative transport. The propagated derivative at a sampled position $t$ is only weighted by the local color difference and the local occupancy respectively (Equation 23). In contrast to exponential or non-exponential (e.g., linear [Vicini et al., 2021a]) volumes, the notion of transmittance disappears.
3.7. Primal rendering
Defining derivative propagation alone is not enough. Many differentiable rendering tasks also need the capability to render the scene to compare to reference images.
However, the transport law in Equation 22 that was designed with derivatives in mind is not suitable for primal rendering. To see why, consider the integrated occupancy along the interval $[0,s]$:
(24) $A=\int_{0}^{s}\alpha^{+}(t)\,\mathrm{d}t$
The value $A$ specifies to what degree the volume masks the mean surface. This masking makes sense when $A\!\in\![0,1]$. But $A$ can be as large as $s$, which potentially exceeds $1$ by a significant amount, and this can cause $L_{\mathrm{i}}^{1}(0)$ to take on unnaturally large or even negative values.
The violation of energy conservation is not surprising and is even desired: many-worlds derivative propagation independently considers the effect of many conflicting worlds, so the question of energy conservation does not even arise. However, naively applying this transport law to normal radiance leads to nonsensical results. To fix this, we scale the line integral by the factor $1/s$, which enforces energy conservation by bounding the effect of the volume so that the most it can do is to completely occludethe base surface.
Following this change, the transport law can also be interpreted as a simple linear blend between foreground and background:
${\hat{L}}_{\mathrm{i}}^{1}(0)$ $=\frac{1}{s}\int_{0}^{s}\alpha^{+}(t)\left[L_{\mathrm{o}}(t)-L_{\mathrm{o}}(s)\right]\,\mathrm{d}t+L_{o}(s)$
(25) $=\frac{1}{s}\int_{0}^{s}\left[\alpha^{+}(t)L_{\mathrm{o}}(t)+(1-\alpha^{+}(t)L_{\mathrm{o}}(s)\right]\,\mathrm{d}t$
Another way to interpret (25) is as an expected value:
(26) ${\hat{L}}_{\mathrm{i}}^{1}(0)=E\left[L_{\mathrm{o}}(X)\right]$
where $X$ is a mixed random variable with definition
(27) $X=\begin{cases}\ t\in(0,s),&\text{with density }\alpha(t)/A,\\ \ s,&\text{with probability }1-A/s,\end{cases}$
The $1/s$ scaling is a natural design choice with a clear interpretation, but it is not unique. Other choices could also be used, as the rendering step is decoupled from the proposed derivative propagation [Nimier-David et al., 2020].
4. Discussion
We are now in a position to examine the proposed method from a theoretical point of view. We make the following observations:
Random surface model
The derivation in Section 3 went through the following sequence of conceptual steps:
  1. (1)
    Perturbing a surface rendering by adding a bit $(\varepsilon)$ of volume.
  2. (2)
    Considering the case $\varepsilon\approx 0$, which removes higher-order effects.
  3. (3)
    Defining the optical properties of this volume.
  4. (4)
    Computing a derivative with respect to $\varepsilon$ to get the final model.
The main ideas in this sequence are (1), (2), and (4), while step (3) was arbitrary. For example, we decided to orient the BRDF of all many-worlds surfaces at $\mathbf{x}$ based on the volume’s mean orientation $\bm{\beta}(\mathbf{x})$, which is simple but also unnecessarily restrictive.
It should be possible to construct versions of this framework with particles following a more general directional distribution. The key insight is to model the perturbation so that worlds become independent from each other, and this would apply here as well (there would just be more worlds to consider at any given point).
Surface derivatives
Prior works on differentiable rendering of surfaces have proposed local formulations [Zhang et al. [2020], Zhang et al. [2023a]] to express derivatives arising from the discontinuous visibility function. These local formulations provide a different viewpoint that can unlock significant efficiency gains. For example, consider a differentiable rendering of a triangle mesh occluding another object. Normally, this would require sampling light paths that graze the silhouette of the occluder, and finding such paths can be computationally expensive. However, it is is equivalent and generally easier to start from the middle by randomly picking a polygonal edge on the occluder along with a tangential direction, and then bidirectionally expanding it into a complete light path. This is referred to as a local formulation because it expresses the gradient computation as a low-dimensional integral over the space where derivatives actually arise (e.g., at silhouette points causing an occlusion).
Recall that many-worlds rendering evaluates a volumetric perturbation of a surface rendering. This is in contrast to prior work that only considers local perturbations of a surface rendering. In the appendix, we start with the classic surface derivatives and analytically integrate them in an extended domain to re-derive the many-worlds derivative transport law (Equation 23).
This derivation grounds the many-worlds method in an established theory and we make the following observations:
  1. (1)
    Our method does the right thing near the surface: the optimization behavior matches surface differentiation strategies without sampling silhouettes or introducing approximations.
  2. (2)
    The extended derivative domain enables materialization of shapes far from the current surface. This is more efficient and robust than locally perturbing surfaces.
  3. (3)
    We unify visibility derivatives and shading derivatives in the extended domain. Prior works separately estimate these two types of derivatives and combine them for backpropagation. The algorithm complexity is greatly reduced with the extended derivative domain.
Figure 6. Importance of the orientation field. We compare reconstructions done with and without an orientation field $\bm{\beta}$. The top row uses an isotropic normal distribution, while the bottom one relies on the anisotropic model introduced in Section 3.4. A lack of orientation information dramatically slows down convergence and produces incorrect meshes.
Sampling strategy
An important detail that must be addressed is how the many-worlds transport equation (Equation 27) should be sampled in a Monte Carlo renderer. Since we don’t know the expected magnitude of $L_{\mathrm{o}}(t)$, the default strategy will simply sample all other terms proportional to their known contribution. In other words, by drawing positions $t$ with density $\alpha(t)/A$ and the surface at $s$ with discrete probability $1-A/s$. While this is a sensible choice for primal rendering, it leads to a chicken-and-egg problem during optimization. It would imply that regions with $\alpha(\mathbf{x})\!\approx\!0$ are essentially never sampled, but we must clearly visit them sometimes to even consider the possibility of placing a surface there.
Recall the many-worlds derivative transport in Eq. 23:
(28) $\partial_{\pi}L_{\mathrm{i}}^{1}(0)=\int_{0}^{s}\big{(}\underbracket{\partial_{\pi}\alpha^{+}(t)\left[L_{\mathrm{o}}(t)-L_{\mathrm{o}}(s)\right]}_{\text{(i)}}+\underbracket{\alpha^{+}(t)\partial_{\pi}L_{\mathrm{o}}(t)}_{\text{(ii)}}\big{)}\,\mathrm{d}t$
Term (i) of this expression states that the derivative of occupancy along the ray is proportional to $L_{\mathrm{o}}(t)-L_{\mathrm{o}}(s)$, which is an unknown quantity that must be estimated. Given that there are no remaining known factors that could be used to sample $t$, the best strategy for this derivative is to uniformly pick points along the ray, estimate the difference, and then accumulate gradients.
This isn’t a new discovery: Nimier-David et al. [2022] discuss the same issue in exponential volumes and address it by introducing generalizations of volume trackers that sample proportional to pure transmittance $T(t)$ instead of the default extinction-weighted transmittance $\mu_{\mathrm{t}}(t)T(t)$. Many-worlds transport has no transmittance, so the solution ends up being even simpler: a uniform sampling strategy suffices.
5. Implementation Details
5.1. Stochastic surface model
The preceding sections provided a complete transport model but left the details of the occupancy and orientation fields unspecified. We outline one possible way to connect the fields to a stochastic surface model here [Williams and Fitzgibbon, 2006; Miller et al., 2024], noting that this is not our contribution and other formulations may be possible as well.
An implicit representation $\Phi(\mathbf{x})$ of a shape determines whether a position $\mathbf{x}$ is inside ($\Phi(\mathbf{x})<0)$, outside ($\Phi(\mathbf{x})>0)$, or on the boundary ($\Phi(\mathbf{x})=0)$. This geometric classification is independent of optical properties—for example, points $\mathbf{x}$ within a refractive material also count as inside ($\Phi(\mathbf{x})<0$).
In our case, the scene models a distribution of surfaces, which turns $\Phi(\mathbf{x})$ into a random variable. The occupancy $\alpha(\mathbf{x})$ then gives the probability of $\mathbf{x}$ being on or inside an object:
(29) $\alpha(\mathbf{x})=\Pr\{\Phi(\mathbf{x})\leq 0\}.$
We model $\Phi(x)$ as a Gaussian process (GP), which means that pointwise evaluations of the implicit function are normally distributed:
(30) $\Phi(\mathbf{x})\sim\mathcal{N}(\mu(\mathbf{x}),\sigma^{2}(\mathbf{x})),$
which yields an explicit form of the occupancy:
(31) $\alpha(\mathbf{x})=\frac{1}{2}\left[1-\operatorname{erf}\left(\frac{\mu(\mathbf{x})}{\sqrt{2\sigma^{2}(\mathbf{x})}}\right)\right].$
We assume a constant variance $\sigma^{2}(\mathbf{x})=\sigma^{2}$ for all $\mathbf{x}$ in this work.
We also use this representation to assign an orientation to every point based on the expected gradient of the implicit function, i.e.,
$\bm{\beta}(\mathbf{x})=\frac{E\left[\nabla\Phi(\mathbf{x})\right]}{\left\|E\left[\nabla\Phi(\mathbf{x})\right]\right\|}=\frac{\nabla\mu(\mathbf{x})}{\left\|\nabla\mu(\mathbf{x})\right\|}.$
Modeling orientation has a crucial impact on the robustness and speed of optimizations (Figure 6).
A GP also has the property that evaluations $\Phi(\mathbf{x}_{1}),\Phi(\mathbf{x}_{2}),\ldots$ follow a joint multivariate normal distribution. Their auto-correlation is often described using a kernel that depends on distance, e.g.:
(32) $\operatorname*{corr}(\Phi(\mathbf{x}),\Phi(\mathbf{y}))=\exp\left(-\gamma\|\mathbf{x}-\mathbf{y}\|^{2}\right)$
where $\operatorname*{corr}()$ refers to Pearson’s correlation coefficient. This ensures that nearby points become increasingly correlated, making sudden jumps in realizations of $\Phi(\mathbf{x})$ unlikely.
Autocorrelation would be a crucial property if our method involved steps such as sampling concrete world realizations from the random variable $S$, or if it evaluated the same world at multiple different locations, requiring careful modeling of the correlation between them. However, our method does not depend on such steps. Its sole interaction with $S$ is through pointwise evaluations of probabilities and extraction of the mean surface
(33) $\bar{S}=\{\mathbf{x}\in\mathbb{R}^{3}\mid\mu(\mathbf{x})=0\}=\{\mathbf{x}\in\mathbb{R}^{3}\mid\alpha(\mathbf{x})=\nicefrac{{1}}{{2}}\}.$
Spatial correlation is therefore not a detail that must be modeled.
The ray-surface intersection at $s$ (Equation 9) considers this mean surface. Our method evolves a statistical distribution of surfaces and materializes them into a concrete surface once the occupancy raises above the threshold of $\nicefrac{{1}}{{2}}$. The mean surface interacts with incident light but otherwise does not directly participate in the optimization (i.e., it is detached in the underlying gradient computation). Despite this, changes in the underlying fields cause it to steadily evolve throughout the optimization process. Section 4 and the appendix explain why this evolution makes sense, and how it can be rigorously justified in terms of prior work on differentiable rendering of geometry.
5.2. Implementation details
Relative boundary motion
Figure 7. Relative motion in the boundary derivative. We render a diffuse, bumpy torus under a constant environment map with direct illumination. The shading derivative is zero everywhere since the lighting is uniform. To reconstruct the concave regions of the torus, we rely on the differentiation of the shadow ray. The derivative needs to be propagated to the shading point as a result of the relative motion of the silhouette boundary.
Previous analysis typically assumes that only the point of interaction is differentiated. However, this is not sufficient for physically-based rendering. In surface derivative theory, the boundary derivative that models the occlusion effects simultaneously propagates the derivative to the silhouette point and the ray origin of the boundary segment (the shading point).
To understand why this propagation is necessary, consider the ambient occlusion-style shading in Figure 7 of a shape with concave regions that never form an occluding silhouette from any viewpoint. The color at any shading point depends solely on whether sampled shadow rays are occluded; in this case, the color depends only on self-occlusion. By optimizing the position of silhouettes, we can reconstruct most of the shape, but the concave regions do not receive any gradients (as the darker appearance of concave regions is not due to shading).
To properly optimize these regions, the system needs to understand that moving the valley floor further inside is equivalent to moving the valley ridge further outside, resulting in reduced radiance at the shading point.
This concept is not new: reparameterization-based methods account for this in the nested reparameterization step, and silhouette segment sampling methods compute the motion of the silhouette segment as the normal motion of the silhouette point relative to the shading point’s motion. We also model this relative motion in many-worlds derivative transport: any derivative propagated to the occupancy field is also be propagated to occupancy at the shading point.
Pseudocode
We implemented a minimalistic version of our method in Mitsuba 3 [Jakob et al., 2022], discarding unnecessary complexity even at the cost of reduced efficiency or quality. It serves as a benchmark that straightforwardly incorporates many-worlds transport into a typical PBR rendering system. The listing below summarizes the necessary changes to a recursive path tracer. Our implementation furthermore adds standard extensions (path termination, emitter sampling, MIS) that are orthogonal to many-worlds rendering.
{minted} [linenos,fontsize=, escapeinside=——, mathescape=true]python def Li(—$\mathbf{x}$—, —$\bm{\omega}$—): # Pick a segment where intersection happens —$k_{\mathrm{mw}}$— = —$\lfloor k_{\mathrm{max}}$— * rand()$\rfloor$— return Li_k(—$\mathbf{x}$—, —$\bm{\omega}$—, —$k_{\mathrm{mw}}$—, 0) def Li_k(—$\mathbf{x}$—, —$\bm{\omega}$—, —$k_{\mathrm{mw}}$—, —$k$—): if —$k$— ¿ —$k_{\mathrm{max}}$—: # Path length exceeds limit return 0 # Incident radiance from mean surface t = ray_intersect(—${\bar{S}}$—, —$\mathbf{x}$—, —$\bm{\omega}$—) —$\mathbf{x}^{\prime}$— = —$\mathbf{x}$— + t * —$\bm{\omega}$— # Advance to surface —$\bm{\omega}^{\prime}$—, —$w$— = sample_bsdf(—$\mathbf{x}^{\prime}$—, -—$\bm{\omega}$—) # Get direction+weight from BSDF L_bg = Le(—$\mathbf{x}^{\prime}$—,-—$\bm{\omega}$—) + Li_k(—$\mathbf{x}^{\prime}$—, —$\bm{\omega}^{\prime}$—, —$k_{\mathrm{mw}}$—, —$k$— + 1) * —$w$— if —$k$— != —$k_{\mathrm{mw}}$—: # Segment is before/after the sampled segment return L_bg # Incident radiance from sampled surface s = rand() * t # Sample a surface —$\mathbf{x}^{\prime}$— = —$\mathbf{x}$— + s * —$\bm{\omega}$— # Advance to sampled surface —$\bm{\omega}^{\prime}$—, —$w$— = sample_bsdf(—$\mathbf{x}^{\prime}$—, -—$\bm{\omega}$—) # Get direction+weight from BSDF occupancy = —$\alpha$—(—$\mathbf{x}^{\prime}$—) # occupancy at sampled point L_fg = Le(—$\mathbf{x}^{\prime}$—,-—$\bm{\omega}$—) + Li_k(—$\mathbf{x}^{\prime}$—, —$\bm{\omega}^{\prime}$—, —$k_{\mathrm{mw}}$—, —$k$— + 1) * —$w$— return lerp(L_bg, L_fg, occ) * (1 / t) # Equation (23)
Figure 8. Multi-view geometry reconstructions using many-worlds derivative transport. For the Deer scene, all $8$ views are behind the object and we only see the front side in the mirror. For Polyhedra, Neptune and Fertility, the object is inside a spherical or cubic smooth glass container. The material is known during optimization: the Dragon has a rough gold material, Polyhedra and Neptune are made of copper oxide, and others are diffuse.
Occupancy and mean surface
We query $\mu(\mathbf{x})$ from a grid-based texture using cubic interpolants. Alternatively, a neural representation [Müller et al., 2022] could be employed for greater flexibility.
To simulate interactions with the detached mean surface (Eq. 22), we use Marching Cubes [Lorensen and Cline, 1987] to extract a triangle mesh from $\mu(\mathbf{x})$, enabling the use of efficient ray-triangle surface intersection methods. Any other isosurface extraction algorithm could in principle be used, since the extraction step does not need to be differentiable. Our pipeline does not involve sphere tracing, delta tracking or other iterative algorithms.
6. Results
Figure 9. Material optimization. This experiment demonstrates joint optimization of geometry and an albedo texture. Our theory mainly focuses on geometric optimization, but it is also compatible with more general inverse rendering pipelines that furthermore target material and lighting.
Figure 10. Assumption-free optimization. Prior surface optimization methods require an initial guess, which affects the quality of their output. We compare our method to an unbiased surface derivative method (SDF reparameterization [Vicini et al., 2022]) using an initialization with a sphere that is either bigger (middle) or smaller (bottom) than the overall size of the target. The experiment shows that the method struggles to reconstruct the target from the smaller sphere, while carving it out of the bigger sphere seems more reliable. Our method (top) can optimize without such assumptions, i.e., starting from empty space.
Figure 11. The benefit of simultaneously optimizing all positions. We replicate an experiment of Zhang et al. [2023a] with our method to reconstruct a chair from a single reference image. The baseline employs preconditioned gradient descent [Nicolet et al., 2021] to locally deform a triangle mesh. In each iteration, the pixels rendering the legs propagate gradients to the scene, causing their progressive extrusion in a thin region of overlap between tentative object and the chair in the reference image. Derivatives outside of the region of overlap are discarded. Our method observes and uses all gradients starting from the very first iteration, enabling faster convergence.
Prior work on physically-based surface inverse rendering often includes comparisons of forward derivatives to other competing methods or finite difference-based reference derivatives. They reveal the change in rendered pixels when perturbing a single scene parameter. However, such a comparison is neither applicable nor meaningful in our method: first, our method computes a different kind of derivative on a higher-dimensional domain, which therefore cannot be compared to existing methods. Second, the need for such comparisons is motivated by the complexity of formulations that deal with discontinuous visibility, but differentiating a volume is “trivial” and therefore not interesting. Because of this, we demonstrate the correctness and performance of our method solely through end-to-end optimizations.
All results presented in this paper exclusively use many-worlds derivatives and explicitly disable traditional surface derivatives, even for the albedo optimization demonstrated in Figure 9. This choice highlights the correctness and efficiency of our method.
Multi-view reconstructions
Figure 8 presents multi-view object reconstructions of a known material in various settings using our theory. All experiments use the Adam optimizer [Kingma and Ba, 2014]. We found that disabling momentum in the early iterations helped avoid excessive changes while the occupancy field was still very far from convergence. Alternatively, a stochastic gradient descent optimizer produces comparable results. During the reconstruction, each view is rendered at a $512\!\times\!512$ resolution.
The last four rows of Figure 8 show reconstructions involving perfectly specular surfaces. No prior PBR method could handle such scenes: reparameterization-based methods would need to account for the extra distortion produced by specular interactions, while silhouette segment sampling methods would need to find directional emitters through specular chains. Both are complex additional requirements that would be difficult to solve in practice, while the problem simply disappears with the many-worlds formulation.
Albedo reconstruction
This article primarily focuses on geometry, but our method can in principle also be used to optimize materials or lighting. Figure 9 presents experiments where we jointly optimize the geometry and albedo texture of an object. We store this spatially varying albedo on an additional 3D volume that parameterizes the BSDF of the many-worlds representation.
Benefits of assumption-free geometry priors
Figure 10 compares the convergence rate of our method to a technique that evolves an SDF using reparameterizations [Vicini et al., 2022]. These experiments demonstrate that our method requires significantly fewer iterations because new surfaces can be materialized from the very first iteration.
We use $20$ random optimization views, with each view rendered at $256\!\times\!256$ pixels. Both methods utilize a geometry grid resolution of $128^{3}$. The time required to perform one iteration of the optimization for one view is $0.25$ seconds for our method, $0.38$ seconds for the large sphere initialization and $0.32$ seconds for the small sphere initialization. Our method benefits from the efficiency of ray-triangle intersections, whereas the SDF representation relies on a more costly iterative sphere tracing algorithm. These measurements also show how sphere tracing slows down with smaller steps due to complicated geometry near a ray, as evidenced by the slower performance of the larger sphere initialization.
Optimizing all positions at once
Figure 11 replicates the chair reconstruction experiment of a recent work by Zhang et al. [2023a] to demonstrate the benefits of the extended parameter space.
Subtractive changes
Figure 12 shows the optimization states for an initialization with random geometry. We used $16$ optimization views and the same setup as in Figure 8. Since the many-worlds derivative extends the surface derivative domain without approximations, our method naturally address scenarios that surface derivatives can handle—in this case, removing superfluous geometry and deforming the rest to reconstruct the desired object.
Timings.
Table 7 lists the average computation time per gradient step and view for several scenes. These timings were measured on an AMD Ryzen 3970X Linux workstation with an NVIDIA RTX 3090 graphics card, which our implementation uses to accelerate ray tracing. We limited the maximum path depth for every scene to a reasonable value. For certain scenes, we also disabled gradient estimation for some path segments. For example, in the shape reconstructions inside a glass object, the first ray segment (from the camera to the glass interface) and the last ray segment cannot intersect the many-worlds representation.
7. Conclusion
Correct handling of discontinuous visibility during physically-based inversion of geometry has presented a formidable challenge in recent years. Instead of proposing yet another method to solve this challenging problem, many-worlds inverse rendering shows that there are completely different ways to approach it. Using a notion of a volumetric perturbation of a surface, our method successfully synthesizes complex geometries from an initially empty scene.
Our work lays the theoretical foundation and validates this theory with a first implementation. However, the details of this implementation are still far from optimal and could benefit from various enhancements. For example, we derive the local orientation from the mean field that also governs occupancy, which is straightforward but also introduces a difficult-to-optimize nonlinear coupling. Extending the model with a distribution of orientations would further allow it to consider multiple conflicting explanations at every point.
Reconstructing an object involves a balance between exploration to consider alternative explanations and exploitation to refine their parameters. Our method aggressively pursues the exploration phase using a uniform sampling strategy but lacks a mechanism to effectively exploit its knowledge. We find that it often reconstructs a good approximation of a complex shape in as little as 20 iterations, impossibly fast compared to existing methods, but then requires 500 iterations for the seemingly trivial task of smoothing out little kinks.
Our implementation of the many-worlds derivative is based on a standard physically-based path tracer, but previous works on differentiable rendering have shown the benefits of moving derivative computation into a separate phase using a local formulation. This phase starts the Monte Carlo sampling process where derivatives locally emerge (e.g., at edges of a triangle mesh in the case of visibility discontinuities). Integrating the local form of our model with occupancy-based sampling could refine the mean surface with a more targeted optimization of its close neighborhood.
The parts of the many-worlds representation beneath the mean surface effectively disappear from view—while we can still locally adjust the surface to reduce its thickness, our method does not consider more drastic changes like punching a hole [Mehta et al., 2023]. Exploring such subtractive perturbations in many-worlds rendering presents an interesting direction for future work.
Figure 12. Subtractive changes. Many-worlds derivatives match surface derivatives when sampled close to the surface. We initialize the scene with dense geometry to demonstrate the robustness of our method against subtractive changes.
Table 1. We measure the average time needed to optimize one $512^{2}$ pixel image for different scenes. The reported time covers all overheads including the primal rendering pass, the uncorrelated derivative propagation pass, the optimizer step and scene update.
Path
AD
spp grad spp time (s)
Heptoroid 2 2 128 32 0.41
Dragon 2 2 128 32 0.42
Deer 4 2 128 32 0.75
Polyhedra 4 2 64 32 1.10
Neptune 5 3 256 32 2.56
Fertility 5 3 256 32 2.06
Acknowledgements.
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No 948846).